### v1

# Non-simple abelian varieties in a family and countable intersections on moduli space

Abstract.
Inspired by the work of Ellenberg, Elsholtz, Hall, and Kowalski, we investigate how the property of the generic fiber of a one-parameter family of abelian varieties being geometrically simple extends to other fibers. In [EEHK09], the authors studied a special case involving specific one-parameter families of Jacobians of curves using analytic methods. We generalize their results, particularly Theorem B, to all families of abelian varieties with big geometric monodromy, employing an arithmetic approach. Our method applies Heath-Brown-type bounds on certain covers with level structures and optimizes the covers to derive the desired results.
1. Introduction
let $\mathcal{A}_{\_}{g}$ be the moduli space of principally polarized abelian varieties of dimension $g$. Let $K$ be a number field and Let $X\in\mathcal{A}_{\_}g/K$ be a rational curve parameterizing a one-dimensional family of abelian varieties. This is equivalent to saying $X\simeq\mathbb{P}^{1}$. Let $A_{\_}{x}$ denote the fiber of this family over $x\in X$ and let $A_{\_}\eta$ be the generic fiber. Assume that the geometric monodromy of $A_{\_}\eta$ modulo $\ell$ is the full symplectic group $\operatorname{Sp}_{\_}2g(\mathbb{F}_{\_}\ell)$ for almost all primes $\ell$. In particular, this implies that $A_{\_}\eta$ is geometrically simple. We are interested in understanding how this property extends to other fibers. To be precise, we ask the following question:
Question 1. What can we say about the number of $x\in X(\mathbb{Q})$ of bounded height such that $A_{\_}{x}$ is geometrically non-simple?
For $t\in X(K)$, let $H(t)$ denote the multiplicative projective height of $t$. Define $S(B)$ as the set
[…]
One may ask whether $S(B)$ is finite. If so, can we obtain an upper bound in terms of $B$? The finiteness of $S(B)$ was proven in [EEHK09, Thm. A] under certain assumptions ensuring large monodromy for the family. Specifically, these assumptions concern a particular type of family of abelian varieties, namely the family
$A_{\_}{f}\to\mathbb{A}^{1}$
of Jacobians of the hyperelliptic curves defined by affine equations
$y^{2}=f(x)(x-t)\text{, t $\in$ $\mathbb{A}^{1}$,}$
for some fixed square-free polynomial $f\in\mathbb{Z}[X]$ of degree $2g$, $g\geq 1$.
The authors of [EEHK09] also provided analytic methods, based on sieve theory, to approach upper bounds for the number of non-simple varieties with height at most $B$.
Theorem 1.1.
[EEHK09, Thm. B] There exist constants $C\geqslant 0$ and $D\geqslant 1$, independent of $f$, such that we have
$|S(B)|\leqslant C\left(g^{2}D(\log 2B)\right)^{11g^{2}}$
for all $B\geqslant 1$.
However, the constant $C$ depends on the largest prime dividing the discriminant of $f$. In this article, we aim to generalize these results to arbitrary rational one-dimensional families of abelian varieties over a number field, without knowing the explicit defining equation of the family. Notably, compared to [EEHK09, Theorem B], the logarithmic terms in our theorem do not depend on $g$ or $K$. To state our theorem, we need the following definitions.
Definition 1.2.
Let $n$ be the transcendence degree of $\mathcal{A}_{g}$ over $K$ and let $N=n+1$. Define $\iota:X\to\mathbb{P}^{N}$ as the restriction of the Hodge embedding of $\mathcal{A}_{g}$ to $X$ (see also Section 2.1 for details).
Therefore, specifying $iota$ is the same as specifying the family $X$.
In the definition of $S(B)$, $H(t)$ is the height of $t$ as an element of $K$. Let $P_{\_}t=\iota(t).$ We now introduce the following definition.
Definition 1.3.
Define $S_{\_}X(B)$ as the set
[…]
Our main theorem is as follows.
Theorem 1.4.
Let $K$ be a number field of degree $d_{\_}K$ over $\mathbb{Q}$. Let $X\in\mathcal{A}_{g}/K$ be a rational curve parameterizing a one-dimensional family of abelian varieties. Let $A_{\_}\eta$ be the generic fiber of this family, and assume that the geometric monodromy of $A_{\_}\eta$ modulo $\ell$ is the full symplectic group $\operatorname{Sp}_{\_}{2g}(\mathbb{F}_{\_}\ell)$ for almost all primes $\ell$. Let $\iota$ be the embedding defined in 1.2 and let $d$ denote the degree of $X$ with respect to $\iota$. Then there exists an absolute constant $C_{\_}{K,g}$, depending on $K$ and $g$, such that
$S_{\_}X(B)\leq C_{\_}{K,g}d^{2}(\log B)^{2+\kappa}$
for sufficiently large $B$.
If we change the definition of the height from $H(P_{\_}t)$ to $H(t)$ and assume that $H(t)\leq B$, then there exists a constant $C_{\_}{\iota}$ depends on $d_{\_}K$, $g$ and the family $X$, and an absolute constant $\kappa$ such that
$|S(B)|\leq C_{\_}{K,g,\iota}d^{2}(\log B)^{2+\kappa}.$
Remark 1.5.
The first part of Theorem 1.4 is uniform, as the method does not require the curve $X$ to have a specific form or be defined by specific equations. The approach also makes it easier to work in generality than in [EEHK09].
If the constant is allowed to depend on $\iota$, then $S(B)$ is bounded above by a constant $C_{\_}\iota$, as demonstrated in the first part of [EEHK09].
The theorem shows that most of the fibers carry the generic property of being simple, while the non-simple fibers exhibit subpolynomial growth with respect to height. However, considering $S(B)$ as the intersection of a rational curve with the moduli space _$\mathcal{M}_{\_}g$ of genus-$g$ curves over the complex numbers, the locus in _$\mathcal{M}_{\_}g$ corresponding to curves with non-simple Jacobians is a countable union of lower-dimensional algebraic subvarieties. Thus, our result implies that most intersection points are irrational.
The strategy for proving Theorem 1.4 is totally different from [EEHK09]. While their approach relies on analytic number theory, our methods are arithmetic in nature. Specifically, we apply Heath-Brown-type bounds on certain covers with level structure and optimize these covers. This approach, used in previous work by the author [Fu] and Ellenberg-Lawrence-Venkatesh [ELV], remains underexplored. We hope this paper highlights the strength of this method, demonstrating its applicability to similar problems.
1.1. Paper outline
In Section 2, we summarize basic results on heights of abelian varieties, height bounds on isogenous abelian varieties, and sharp bounds on rational points of bounded height on curves by Binyamini-Cluckers-Kato. In Section 3, we construct congruence covers to capture rational points on $X$ parameterizing non-simple abelian varieties and provide a moduli description of these covers. In Section 4, we apply Faltings’ theorem on the heights of isogenous abelian varieties to derive an upper bound for height changes between rational points on $X$ and their lifts to the covers. In Section 5, we combine these results with the theorem of Binyamini-Cluckers-Kato to optimize the covers and complete the proof of Theorem 1.4.
1.2. Acknowledgement
The author thanks Jordan Ellenberg for many helpful suggestions and discussions.
2. Preliminaries
2.1. Heights on abelian varieties
Let $K$ be a number field. The graded $K$-algebra of modular forms on $\mathcal{A}_{g}$ defined over $K$ is finitely generated. Moreover, there exists a positive weight $w$ such that modular forms of weight $w$ defined over $K$ realize a projective embedding of $\mathcal{A}_{g}$ and every element of the function field $K(\mathcal{A}_{g})$ can be expressed as a quotient of two modular forms of the same weight defined over $K$.
In fact, letting $\mathcal{A}^{\operatorname{univ}}\to\mathcal{A}_{\_}g$ denote the universal abelian variety, one can define an ample line bundle _$\mathcal{L}_{\_}g:=\operatorname{det}\mathrm{Fil}^{1}H_{\_}{dR}^{1}\left(\mathcal{A}^{\operatorname{univ}}/\mathcal{A}_{g}\right)$ on $\mathcal{A}_{g}$, referred to as the Hodge bundle. For $w\geq 1$, the algebraic sections of _$\mathcal{L}_{\_}g^{\otimes w}$ are exactly the modular forms of weight $w$ on $\mathcal{A}_{g}$(cf. [Kie20, Theorem 2.5].) For instance, as shown by Igusa [Igu] and Streng [Streng], the graded $\mathbb{Q}$-algebra of modular forms on $\mathcal{A}_{\_}2$ over $\mathbb{Q}$ is generated by $I_{\_}4,I_{\_}6^{\prime},I_{\_}{10}$ and $I_{\_}{12}$ of respective weights 4, 6, 10 and 12. The function field is generated by the Igusa invariants:
$j_{\_}1=\frac{I_{\_}4I_{\_}6^{\prime}}{I_{\_}{10}},\quad j_{\_}2=\frac{I_{\_}4^{2}I_{\_}{12}}{I_{\_}{10}^{2}},\quad j_{\_}3=\frac{I_{\_}4^{5}}{I_{\_}{10}^{2}}.$
Thus, the embedding of $\mathcal{A}_{\_}2$ is defined by modular forms of weight 20.
There are various types of heights can be defined for an abelian variety $A\in\mathcal{A}_{g}$ over $\overline{\mathbb{Q}}$. One can defined the Faltings height $h_{\_}F(A)$ in terms of the Arakelov degrees of the Hodge bundle _$\mathcal{L}_{\_}g$. Alternatively, one may define the $j$-height of $A$ as the projective height of the generators $j_{\_}1,\dots,j_{\_}{n}$ of $\mathbb{Q}(\mathcal{A}_{g})$, where $n$ is the transcendence degree over $\mathbb{Q}$ and $j_{\_}{n+1}$ generates the remaining finite extension
$h_{\_}j(A)=h(j_{\_}1(A);\dots;j_{\_}{n+1}(A);1).$
2.2. Height bounds on isogenous abelian varieties
Let $A^{\prime}$ be an abelain variety such that there exists an isogeny $\varphi:A\to A^{\prime}$. We want to relate the height of $A^{\prime}$ to that of $A$ under cartain conditions related to the geometry of the moduli space. For the Faltings height, we have the following theorem by Faltings [Faltings1983, Lem. 5]
Proposition 2.1.
$|h\left(A^{\prime}\right)-h\left(A\right)|\leq\frac{1}{2}\log(\operatorname{deg}(\phi)).$
There are results of a similar flavor for the $j$-heights of isogenous abelian varieties. In particular, we point out the result by Kieffer in [Kie20, Cor. 5.10] and note that the choice of the type of height does not make a significant difference to our final result. This is due to the existence of nice comparison theorems between different types of heights, and the role of the optimization process in our approach. See Section 5 for further details.
2.3. Uniform bound for rational points of bounded height on curves and dimension growth
We shall make heavy use of the following theorem of Binyamini-Cluckers-Kato [BCK24].
Let $K$ be a global field. Suppose that $K$ is of degree $d_{\_}K$ over $\mathbb{Q}$ or over $\mathbb{F}_{\_}q(t)$ and in the latter case that $K$ is separable over $\mathbb{F}_{\_}q(t)$ with $\mathbb{F}q$ as its field of constants, where $q$ is a power of a prime number. Let $C\subset\mathbb{P}_{\_}{\mathbb{Q}}^{n}$ be an irreducible algebraic curve of degree $d$. Finally let $C(K,H)$ denote the set of $K$-rational points on $C$ with absolute projective multiplicative height at most $H$.
Theorem 2.2.
[BCK24, Thm. 1] Let $n>1$ be given. There is a constant $c=c(K,n)$ and an absolute constant $\kappa$ such that for any $d>0$, any irreducible algebraic curve $C$ of degree $d$ in _$\mathbb{P}_{\_}K^{n}$ and any $H>2$ one has
(2.1) #$\#C(K,H)\leqslant cd^{2}H^{\frac{2d_{\_}K}{d}}(\log H)^{\kappa}.$
The quadratic dependence in (2.1) is optimal by [CCDN19, Sec. 6], from which one can finds a constant $c^{\prime}=c^{\prime}(K)>0$, arbitrarily large values $d,H$ and irreducible curves $C$ in _$\mathbb{P}_{\_}{\mathbb{Q}}^{2}$ of degree $d$ which witness the lower bound
$c^{\prime}d^{2}H^{\frac{2d_{\_}K}{d}}\leqslant\#C(K,H).$
2.4. Parabolic subgroups of the symplectic groups
Let $G=\operatorname{Sp}(\psi)$ be the symplectic group for a symplectic space $(V,\psi)$ of dimension $2n$, where $n$ is a positive integer and $k$ is a field. We call an increasing flag $F={F^{i}}_{\_}i$ of nonzero proper subspaces of $V$ isotropic if each $F^{i}$ is isotropic with respect to the symplectic form $\psi$. Each parabolic $k$-subgroup of $G$ is the stabilizer of some isotropic flag. In particular, a maximal parabolic subgroup of $G$ is exactly the stabilizer of a minimal non-empty isotropic flag; i.e., the $G$-stabilizer of a non-zero isotropic subspace $F^{1}\subset V$. It is a standard fact that there are $n$ such conjugacy classes, corresponding to a choice of vertex to remove from the Dynkin diagram and to the dimension of $F^{1}$, which lies between $1$ and $n$.
3. Construction and integrality of congruence covers
3.1. The congruent covers
We study level structures on $\mathcal{A}_{\_}g$, the moduli space of principally polarized abelian varieties, formed by certain congruence subgroups $\Gamma$ of $\Gamma(1):=\mathrm{Sp}_{\_}{2g}(\mathbb{Z})$. We are particularly interested in the standard level subgroups $\Gamma(\ell)$, which encode principally polarized abelian varieties $A$ equipped with a symplectic basis of $A[\ell]$. This level structure is denoted by $\mathcal{A}{g,\Gamma(\ell)}$. The cover $\mathcal{A}_{\_}{g,\Gamma(\ell)}$ carries an action of $\mathrm{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z}).$
Let $H$ be a subgroup of $\mathrm{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ that contains a maximal parabolic subgroup of $\mathrm{Sp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z}).$ There are several ways to represent the congruence quotient $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$: either via the cover $q:\mathcal{A}{g,\Gamma(\ell)}/H\to\mathcal{A}_{\_}g$, or via the modular correspondence in $\mathcal{A}_{\_}g\times\mathcal{A}g$, where a point on $\mathcal{A}{g,\Gamma(\ell)}/H$ can be identified with a pair $(A,F^{1})$ with
$F^{1}\subset A[\ell]\simeq(\mathbb{Z}/\ell\mathbb{Z})^{2g}.$
Since $F^{1}$ is isotropic, the pair can also be expressed as $(A,B)$, where $B:=A/F^{1}$ is the quotient abelian variety, which is principally polarized and isogenous to $A$ via an isogeny $\varphi:A\to B$.
Our goal is to bound the number of points with bounded height of $X(K)$, such that the corresponding moduli points are non-simple. Let $\mathcal{A}_{\_}x$ be the fiber over $x\in X(K)$. If $\mathcal{A}_{\_}x$ is non-simple, then the image of the Galois representation $\operatorname{Gal}_{\_}K\to\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ is conjugate to some block matrices and therefore its intersection with $\operatorname{Sp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$ is contained in some maximal parabolic subgroups $H$. Let $\tilde{X}_{\_}H\subset\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ be the lift of $X$ and $\pi:\tilde{X}_{\_}H\to X$ be the restriction of $q$ to $\tilde{X}_{\_}H$. Our task now is to bound the number of points in $X(K)$ that lie in $\pi(\tilde{X}_{\_}H(K))$. The modular interpretation enables us to construct suitable projective embeddings to control the change of heights when lifting the base curve $X$. We will discuss these details in Section 4. We conclude this subsection with a summary in the following corollary:
Corollary 3.1.
Every $k$-rational point $t\in S(B)$ lifts to some congruence cover $\mathcal{A}_{\_}{g,\Gamma(\ell)}/H$ for some prime integer $\ell$ and maximal parabolic subgroup $H\leq\operatorname{GSp}_{\_}{2g}(\mathbb{Z}/\ell\mathbb{Z})$.
3.2. The lift is integral
Let $U\subset X$ be the smooth locus of the family, and $\eta$ denote the generic point. Let $\ell$ be a prime integer. We say a family $\mathcal{A}$ of abelian varieties over $U$ has big monodromy modulo $\ell$ if the image of the geometric monodromy of $A_{\_}\eta$
$\varphi:\pi_{\_}1(U)\to\operatorname{Aut}(\mathcal{A}_{\_}{\eta}[\ell])$
is the full symplectic group $\operatorname{Sp}_{\_}{2g}.$ Similarly, we say an abelian variety $A$ over a field $k$ has big monodromy modulo $\ell$ if the image of Galois representation on $\ell$-torsion points contains the full symplectic group. We say the family, or an abelian variety, has big monodromy if it has big monodromy modulo $\ell$ for almost all primes $\ell$.
For any $g\geq 1$, there exists a constant _$\ell_{\_}1(g)\geq 1$ such that if _$\ell\geq\ell_{\_}1(g)$ and $A$ is an abelian variety of dimension $g$ over a field $k$ with big monodromy modulo $\ell$, then $\operatorname{End}(A)=\mathbb{Z}$ and, in particular, $A$ is geometrically simple ([EEHK09, Prop. 4]). However, the converse is not always true. For example, let $\mathcal{A}\to C$ be a family of abelian surfaces with real multiplication by $\mathbb{Q}(\sqrt{d})$. The monodromy group of $A_{\_}\eta$ modulo $\ell$ is either $\operatorname{GL}_{\_}2(\mathbb{F}_{\_}\ell)\times\operatorname{GL}_{\_}2(\mathbb{F}_{\_}\ell)$ or $\operatorname{GL}(\mathbb{F}_{\_}\ell^{2})$, yet $A_{\_}\eta$ can still be geometrically simple. To prove the integrality, we assume that the family $\mathcal{A}$ has big monodromy.
Proposition 3.2.
Assuming that the family $\mathcal{A}$ over $X$ has big monodromy, the cover $\tilde{X}_{\_}H\to X$ is an integral curve.
Proof.
When $H$ has full determinant, the Weil pairing issue is resolved. For sufficiently large $\ell$, the Galois image of the $-\ell$ geometric monodromy representation contains the full symplectic group. Consequently, the monodromy of $\pi_{\_}1(U)$ acts transitively on the right cosets $\operatorname{Sp}_{\_}{2g}/H$, implying that the curve $\tilde{X}_{\_}H$ is connected. ∎
The big monodromy condition is known for a large class of abelian varieties defined over finitely generated fields of arbitrary characteristic. The most prominent result of this type is the classical theorem of Serre (cf. [Ser84, 2.2.7], [Ser85, Lemme 1]) for the number field case. Its generalization to finitely generated fields of characteristic zero is well-known: If $A$ is an abelian variety over a finitely generated field $k$ of characteristic zero, with $\operatorname{End}_{\_}{\bar{k}}(A)=\mathbb{Z}$ and $\operatorname{dim}(A)=2,6$ or odd, then $A/k$ has big monodromy.
Let $A$ be an abelian variety over a finitely generated field $K$. We say $A$ is of Hall type if $\operatorname{End}(A)=\mathbb{Z}$ and $A$ has semistable reduction of toric dimension one at a place of the base field $K$. In the case where $K$ is a global field, Hall showed that all abelian varieties of Hall type have big monodromy (cf. [Hal08], [Hal11]). Later in 2013, Arias-de-Reyna, Wojciech Gajda and Sebastian Petersen extended the result to all dimensions and for arbitrary finitely generated infinite ground field $K$ in [AGP13].
In combination with the previous proposition we obtain the following corollary
Corollary 3.3.
Let $\mathcal{A}$ be an abelian variety over $K=k(X)$. Assume that conditions $i)$ or $ii)$ are satisfied.
  • i)
    $\mathcal{A}$ is Hall type.
  • ii)
    $\operatorname{char}(K)=0,\operatorname{End}(A)=\mathbb{Z}$ and $\operatorname{dim}(A)=2,6$ or odd.
Then $\tilde{X}_{\_}H\to X$ is integral.
4. Bounding change of heights
In this section, we provide an upper bound on the height of rational points in $\tilde{X}_{\_}{H}(K)$ lying over a point $P_{\_}t\in X(K)$, in terms of the height of $t$ and the level $\ell$. The main result of this section is Proposition 4.1.
Let $\iota_{\_}H^{\prime}$ denote the restriction of the product of Hodge embeddings to $\tilde{X}_{\_}{H}$ and let $\iota_{\_}H$ be the composition of $\iota_{\_}H^{\prime}$ with the Segre embedding. Corollary 3.1 asserts that, for a suitable choice of $\ell\in\mathbb{Z}$, there exist congruence covers of $X$ corresponding to maximal parabolic subgroups, such that a rational point $t\in X(K)$ with height at most $B$ lifts to a rational point on the cover. These covers can be arranged into the following commutative diagram:
[…]
Proposition 4.1.
Fix $\ell\in\mathbb{N}$. Let $t\in K$ be a rational point such that $t\in S(B)$ for some $B$. Let $P_{\_}t$ denote the point on $X$ parameterizing the abelian variety $A_{\_}t$, and let $\tilde{P_{\_}t}$ be a rational lift of $P_{\_}t$ to $\tilde{X}_{\_}{H}(K)$, parameterizing a pair of abelian varieties $(A_{\_}t,A^{\prime}_{\_}t)$ such that $A_{\_}t$ and $A^{\prime}_{\_}t$ are linked by an isogeny of degree $\ell^{g}$. Let $H(\iota_{\_}{H}(\tilde{P_{\_}{t}}))$ be the projective multiplicative height of $\tilde{P_{\_}{t}}$ with respect to $\iota_{\_}{H}$. Let $d$ be the degree of $\iota$. Then
$H(\iota_{\_}{H}(\tilde{P_{\_}{t}}))\leq C_{\_}\iota\ell^{\frac{g}{2}}B^{2d}$
where $C_{\_}\iota$ is a constant depending on $j_{\_}1,\dots,j_{\_}{n+1}$.
Proof.
From [Sil09, VIII, Theorem 5.6], we know that
$H(\iota_{\_}{H}(P_{\_}t))=H(A)\leq C_{\_}\iota B^{d}.$
The constant $C_{\_}\iota$ is computable if the coefficient of $j_{\_}1,\dots,j_{\_}{n+1}$ are known.
We deduce from the fact that the height of the product of two homogeneous polynomials in a set of variables is the product of their heights:
$H(\iota_{\_}{H}(\tilde{P_{\_}{t}}))=H(A)H(A^{\prime}).$
Combining this with Proposition 2.1, the proposition follows. ∎
5. Proof of the main theorem
We are now ready to prove Theorem 1.4. We begin with the following lemma, where $H$ is fixed to be the stabilizer of a maximal isotropic subspace.
Lemma 5.1.
We have
$\deg(\iota_{\_}H)=d\binom{2N}{N}\prod_{\_}{i=1}^{g}(\ell^{i}+1)$
Proof.
The degree of $\varphi^{*}\mathcal{L}_{\_}g|_{\_}{\tilde{X}_{\_}H}$ is $\deg(\varphi)\cdot\deg(\mathcal{L}_{\_}g|_{\_}{X})$ and it is not hard to see that
$\deg(\varphi)=|\operatorname{Sp}_{\_}{2g}(\mathbb{F}_{\_}\ell)/H|=\prod_{\_}{i=1}^{g}(\ell^{i}+1).$
The degree of the Segre variety is a standard fact. ∎
Proof of Theorem 1.4.
First, we prove the second part of the theorem. Let $S_{\_}{B,\ell,H}$ denote the set of rational points on $\tilde{X}_{\_}{H}(K)$ that are liftings of $P_{\_}{t}$ for some $t\in S(B)$. Using Theorem 2.2, Proposition 4.1 and Lemma 5.1, we deduce:
(5.1) $S_{\_}{B,\ell,H}$ $\leq C\deg(\iota_{\_}H)^{2}(H(\iota_{\_}{H}(\tilde{P_{\_}{t}})))^{\frac{2d_{\_}K}{\deg(\iota_{\_}H)}}(\log H(\iota_{\_}{H}(\tilde{P_{\_}{t}})))^{\kappa}$
(5.2) $\leq C_{\_}{K,g}d^{2}\prod_{\_}{i=1}^{g}(\ell^{i}+1)^{2}(\ell^{\frac{g}{2d}}B)^{\frac{C^{\prime}_{\_}{g,K,\iota}}{\prod_{\_}{i=1}^{g}(\ell^{i}+1)}}(\log C_{\_}{\iota}+\frac{1}{2}\log\ell+2d\log B)^{\kappa}$
We now choose $\ell$ to optimize the power-of-$B$ term, obtaining an upper bound for $S_{\_}{B,\ell,H}$ as a consequence. We claim that $\ell$ can be chosen such that $\ell\sim(\log B)^{\frac{2}{g^{2}+g}}$. Indeed,
$B^{\frac{C^{\prime}_{\_}{g,K,\iota}}{\prod_{\_}{i=1}^{g}(\ell^{i}+1)}}$ $<B^{\frac{C^{\prime}_{\_}{g,K,\iota}}{\ell^{\frac{g^{2}+g}{2}}}}$
$=e^{\frac{C^{\prime}_{\_}{g,K,\iota}}{\ell^{\frac{g^{2}+g}{2}}}\log B}$
$\sim e^{C^{\prime}_{\_}{g,K,\iota}}$
which is bounded absolutely by a constant depends on $g,K$ and $\iota$. Therefore, we see that
$S_{\_}{B,\ell,H}$ $\leq C_{\_}{K,g,\iota}d^{2}(\log B)^{2+\kappa}.$
The bound for $S_{\_}X(B)$ follows from a similar argument as above. Let $S_{\_}{X,B,\ell,H}$ denote the set of rational points on $\tilde{X}_{\_}{H}(K)$ that are liftings of $P_{\_}{t}$ for some $t\in S_{\_}X(B)$. Then we have
$S_{\_}{X,B,\ell,H}\leq C_{\_}{K,g}d^{2}(\log B)^{2+\kappa}$
where $C_{\_}{K,g}$ depends on $K$ and $g$. ∎
Compliance with ethical standards
Conflicts of interest
The author states that there are no conflicts of interest.
Data availability
Data sharing not applicable to this article as no data sets were generated or analyzed.